# Summaries as Centroids for Interpretable and Scalable Text Clustering

### v4

@@ 5.1 Results with static data
Comparing our approaches with different LLMs and state‑of‑the‑art LLM‑based clustering methods. Table 3 reports the number of LLM calls (prompts) and mean ACC/NMI for k-NLPmeans (LSA-multiple) and k-LLMmeans (FS-multiple) on three benchmarks with e5-large embeddings. For k-LLMmeans we vary the LLM generator, including GPT-3.5, and observe stable performance across LLMs, indicating robustness to the specific model. By design, k-NLPmeans uses zero LLM calls. For context, we also include results from Feng et al. [20]: ClusterLLM (fine-tuned) and IDAS use GPT-3.5 with the same e5-large embeddings, while LLMEdgeRefine uses the stronger Instructor embeddings [29]. Our GPT-3.5 configuration of k-LLMmeans achieves comparable, sometimes slightly lower ACC/NMI, but with much fewer LLM calls (independent of dataset size) and no fine-tuning. k-NLPmeans trails k-LLMmeans slightly but remains competitive without any LLM usage. We also evaluate BERTopic and apply our variants to its embeddings. In this setting, our methods achieve results extremely close to the best LLM-based models while requiring none or only a small number of LLM calls. Overall, our framework offers a favorable quality–cost trade-off: k-NLPmeans for zero-inference-cost deployments, and k-LLMmeans for higher accuracy under a small, fixed summarization budget.
@@ 6 Sequential (mini-batch) experiments
We use yearly posts from 2020 to 2023 from 35 StackExchange sites [60]. Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods. We release this dataset alongside our submission (see Appendix A.2 for details). For each yearly subset, we split the data into $b=\lceil\tfrac{n}{10000}\rceil$ equal-sized batches $D_{1},\dots,D_{b}$} in chronological order, where $n$ is the number of documents for that year. We run the mini-batch k-LLMmeans algorithm in its four variants described in Section 5 (for practical reasons we set $m=50$ for the full cluster variant). We compare with three baselines: mini-batch k-means with standard random sampling, sequential mini-batch k-means with $b$ chronological batches, and standard k-means on the full dataset. We use ground-truth clusters, text-embedding-3-small for embeddings, GPT-4o for the summarization step, and five different seeds.
@@ 7 Case study
To demonstrate the interpretability of our method in capturing the evolution of clusters within sequential data, we present a case study using posts from the AI site in the 2021 Stack Exchange dataset [60]. We apply our mini-batch k-LLMmeans algorithm with three equal-length batches and a total of ten clusters. We use the instruction “The following is a cluster of questions from the AI community. Write a single question that represents the cluster”.
Interpretation. Our interpretable clustering enables temporal analysis of topic change, since each centroid is a human-readable summary rather than a latent vector. Figure 2 shows this for three major clusters: Image Model Optimization, AI Evolution and Challenges, and Advanced NLP Techniques. Across 2021, these clusters drift from fundamentals toward integration-at-scale, mirroring how tools, models, and deployment pressures matured. Image Model Optimization begins with robustness anxieties (adversaries, occlusion, scale) around mainstream detectors/segmenters, pivots mid-year to “how do I implement this well?” (transfer learning, augmentation, preprocessing), and ends with production hardening (class imbalance, specialized losses, edge efficiency) as strong pretrained models and libraries made setup routine and pushed bottlenecks to data and deployment [8, 32]. AI Evolution and Challenges moves from historical “what is AI becoming?” to curiosity about large generative/multimodal systems, probably sparked by public releases like CLIP/DALL·E and their failure modes, and finishes with system-level synthesis (neuro-symbolic integration, RL control, compute efficiency, safety) as enterprise adoption and policy attention grow [49, 51, 40, 42, 19]. Advanced NLP Techniques tracks the transformer boom: early questions focus on which models and preprocessing to use; mid-year attention shifts to measuring semantics beyond BLEU (e.g., BERTScore, BLEURT) as long-context and domain-shift issues surface; late-year concerns are product-driven, sequence limits, domain vocab, multilingual/cross-modal alignment; reflecting rapid uptake of pretrained encoders/decoders in industry [10, 72, 57, 68]. Practically, these trajectories can power better post categorization, searchability, answer routing, and trend detection. More broadly, they showcase how our mini-batch k-LLMmeans exposes readable centroids enabling end-to-end interpretability for sequential text streams, letting practitioners track topic evolution, attribute causes, and make transparent, timely decisions in dynamic corpora.
Figure 2: Sequential evolution of the LLM-generated centroids for three primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site [60]. Main aspects are manually highlighted.
@@ 8 Related work
Traditional clustering. Hierarchical methods [30, 6] build tree-structured representations of nested document relationships. Density-based approaches like DBSCAN [18] and graph-based methods detect clusters of arbitrary shapes, while spectral clustering [41] leverages eigen-decomposition to uncover complex structures. Model-based techniques; including Gaussian mixture models [15] and recent neural network frameworks [74, 25, 69, 70, 67]; provide probabilistic clustering formulations. Additionally, topic modeling methods, from probabilistic latent semantic analysis [24] to latent Dirichlet allocation [7], capture word co-occurrence patterns and latent topics. BERTopic [23] offers an embedding–reduction–density pipeline for interpretable topics, while DeepCluster [11] uses clustering to iteratively train encoders in a self-supervised manner (orthogonal to our fixed-embedding setting).
LLM-based clustering. Viswanathan et al. [64] employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering. Zhang et al. [73] propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity. Complementary approaches [63, 47] show that context-derived representations capture subtle semantic nuances beyond traditional embeddings. Wang et al. [66] introduce a goal-driven, explainable method that employs natural language descriptions to clarify cluster boundaries, while [13] present IDAS for intent discovery using abstractive summarization. Feng et al. [20] propose LLMEdgeRefine an iterative mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness.
@@ 9 Discussion
We release anonymized code to reproduce all results in this paper at https://anonymous.4open.science/r/summaryCentroids-877C/ (see the README for instructions).
@@ A.1 Non-streaming datasets
We evaluate our clustering approach on four benchmark datasets. Bank77 [12]: Consists of 3,080 customer queries related to banking services, categorized into 77 distinct intents. CLINC [33]: A diverse set of 4,500 queries spanning 150 intent classes across multiple domains, designed for open-domain intent classification. GoEmo [16]: Contains 2,984 social media posts annotated with 27 fine-grained emotion categories. We removed the neutral expressions to address data imbalance and retained only entries with a single, unique emotion. MASSIVE [21]: Comprises 2,974 English-language virtual assistant utterances grouped into 18 domains and 59 intent categories.
@@ A.2 New compiled dataset for testing text-Streaming Clustering Algorithms
We extract and unify a challenging data stream comprising unique archive posts collected from 84 Stack Exchange sites [60]. Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods. Our raw dataset spans 84 domains, each containing at least 20 posts per year from 2018 to 2023 (with post lengths ranging from 20 to 1000 characters), totaling 499,359 posts. For our experiments, we focus on posts from 2020 to 2023 and further filter out labels that do not exceed 500 posts in 2023. The resulting subset comprises 35 distinct groups and 69,147 posts. Both the raw and clean data are provided with this paper. Stack Exchange content is licensed under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) license.
@@ Appendix D Algorithms
    $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\texttt{k-means++}(\{d_{1},\dots,d_{n}\},k)$;
@@
       $\boldsymbol{\mu}_{j}\leftarrow\text{Embedding}\left(\text{j}^{\text{th}}\text{ cluster summary}\right)$;
@@
       $\boldsymbol{\mu}_{j}\leftarrow\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}$;
@@
return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{s_{1},\dots,s_{k}\}$
@@
$\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\}\leftarrow\{{\bf 0},\dots,{\bf 0}\}$;
@@
 $\{\boldsymbol{\mu}^{*}_{1},\dots,\boldsymbol{\mu}^{*}_{k}\},\{C^{*}_{1},\dots,C^{*}_{k}\},S_{b}\leftarrow\texttt{k-NLPmeans}(D_{i},k,I,m,l,T)\text{ or }\texttt{k-LLMmeans}(D_{i},k,I,m,l,T)$;
@@
return $\{\boldsymbol{\mu}_{1},\dots,\boldsymbol{\mu}_{k}\},\{S_{1},\dots,S_{b}\}$

### v5

@@ 5.1 Results with static data
Comparing our approaches with different LLMs and state‑of‑the‑art LLM‑based clustering methods. Table 3 reports the number of LLM calls (prompts) and mean ACC/NMI for k-NLPmeans (LSA-multiple) and k-LLMmeans (FS-multiple) on three benchmarks with e5-large embeddings. For k-LLMmeans we vary the LLM generator, including GPT-3.5, and observe stable performance across LLMs, indicating robustness to the specific model. By design, k-NLPmeans uses zero LLM calls. For context, we also include results from Feng et al. (2024): ClusterLLM (fine-tuned) and IDAS use GPT-3.5 with the same e5-large embeddings, while LLMEdgeRefine uses the stronger Instructor embeddings (Jin et al., 2023). Our GPT-3.5 configuration of k-LLMmeans achieves comparable, sometimes slightly lower ACC/NMI, but with much fewer LLM calls (independent of dataset size) and no fine-tuning. k-NLPmeans trails k-LLMmeans slightly but remains competitive without any LLM usage. We also evaluate BERTopic and apply our variants to its embeddings. In this setting, our methods achieve results extremely close to the best LLM-based models while requiring none or only a small number of LLM calls. Overall, our framework offers a favorable quality–cost trade-off: k-NLPmeans for zero-inference-cost deployments, and k-LLMmeans for higher accuracy under a small, fixed summarization budget.
@@ 6 Sequential (mini-batch) experiments
We use yearly posts from 2020 to 2023 from 35 StackExchange sites (StackExchange, 2024). Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods. We release this dataset alongside our submission (see Appendix A.2 for details). For each yearly subset, we split the data into $b=\lceil\tfrac{n}{10000}\rceil$ equal-sized batches $D_{1},\dots,D_{b}$} in chronological order, where $n$ is the number of documents for that year. We run the mini-batch k-LLMmeans algorithm in its four variants described in Section 5 (for practical reasons we set $m=50$ for the full cluster variant). We compare with three baselines: mini-batch k-means with standard random sampling, sequential mini-batch k-means with $b$ chronological batches, and standard k-means on the full dataset. We use ground-truth clusters, text-embedding-3-small for embeddings, GPT-4o for the summarization step, and five different seeds.
@@ 7 Case study
To demonstrate the interpretability of our method in capturing the evolution of clusters within sequential data, we present a case study using posts from the AI site in the 2021 Stack Exchange dataset (StackExchange, 2024). We apply our mini-batch k-LLMmeans algorithm with three equal-length batches and a total of ten clusters. We use the instruction “The following is a cluster of questions from the AI community. Write a single question that represents the cluster”.
Interpretation. Our interpretable clustering enables temporal analysis of topic change, since each centroid is a human-readable summary rather than a latent vector. Figure 2 shows this for three major clusters: Image Model Optimization, AI Evolution and Challenges, and Advanced NLP Techniques. Across 2021, these clusters drift from fundamentals toward integration-at-scale, mirroring how tools, models, and deployment pressures matured. Image Model Optimization begins with robustness anxieties (adversaries, occlusion, scale) around mainstream detectors/segmenters, pivots mid-year to “how do I implement this well?” (transfer learning, augmentation, preprocessing), and ends with production hardening (class imbalance, specialized losses, edge efficiency) as strong pretrained models and libraries made setup routine and pushed bottlenecks to data and deployment (Bochkovskiy et al., 2020; Kolesnikov et al., 2021). AI Evolution and Challenges moves from historical “what is AI becoming?” to curiosity about large generative/multimodal systems, probably sparked by public releases like CLIP/DALL·E and their failure modes, and finishes with system-level synthesis (neuro-symbolic integration, RL control, compute efficiency, safety) as enterprise adoption and policy attention grow (Radford et al., 2021; Ramesh et al., 2021; Nayak, 2021; OpenAI, 2021; European Commission, 2021). Advanced NLP Techniques tracks the transformer boom: early questions focus on which models and preprocessing to use; mid-year attention shifts to measuring semantics beyond BLEU (e.g., BERTScore, BLEURT) as long-context and domain-shift issues surface; late-year concerns are product-driven, sequence limits, domain vocab, multilingual/cross-modal alignment; reflecting rapid uptake of pretrained encoders/decoders in industry (Brown et al., 2020; Zhang et al., 2020b; Sellam et al., 2020; Xue et al., 2021). Practically, these trajectories can power better post categorization, searchability, answer routing, and trend detection. More broadly, they showcase how our mini-batch k-LLMmeans exposes readable centroids enabling end-to-end interpretability for sequential text streams, letting practitioners track topic evolution, attribute causes, and make transparent, timely decisions in dynamic corpora.
Figure 2: Sequential evolution of the LLM-generated centroids for three primary clusters during the three batches of the sequential mini-batch k-LLMmeans process applied to 2021 posts from the AI Stack Exchange site (StackExchange, 2024). Main aspects are manually highlighted.
@@ 8 Related work
Traditional clustering. Hierarchical methods (Johnson, 1967; Blashfield and Aldenderfer, 1978) build tree-structured representations of nested document relationships. Density-based approaches like DBSCAN (Ester et al., 1996) and graph-based methods detect clusters of arbitrary shapes, while spectral clustering (Ng et al., 2001) leverages eigen-decomposition to uncover complex structures. Model-based techniques; including Gaussian mixture models (Dempster et al., 1977) and recent neural network frameworks (Zhou et al., 2019; Huang et al., 2014; Yang et al., 2016; Zhang et al., 2021; Xie et al., 2016); provide probabilistic clustering formulations. Additionally, topic modeling methods, from probabilistic latent semantic analysis (Hofmann, 2001) to latent Dirichlet allocation (Blei et al., 2003), capture word co-occurrence patterns and latent topics. BERTopic (Grootendorst, 2022) offers an embedding–reduction–density pipeline for interpretable topics, while DeepCluster (Caron et al., 2018) uses clustering to iteratively train encoders in a self-supervised manner (orthogonal to our fixed-embedding setting). Jia and Diaz-Rodriguez (2026) also incorporate modern embeddings into text segmentation within classical change point detection methods.
LLM-based clustering. Viswanathan et al. (2024) employ LLMs to augment document representations, generate pseudo pairwise constraints, and post-correct low-confidence assignments for query-efficient, few-shot semi-supervised clustering. Zhang et al. (2023) propose ClusterLLM, which uses instruction-tuned LLMs via interactive triplet and pairwise feedback to cost-effectively refine clustering granularity. Complementary approaches (Tipirneni et al., 2024; Petukhova et al., 2025) show that context-derived representations capture subtle semantic nuances beyond traditional embeddings. Wang et al. (2023) introduce a goal-driven, explainable method that employs natural language descriptions to clarify cluster boundaries, while (De Raedt et al., 2023) present IDAS for intent discovery using abstractive summarization. Feng et al. (2024) propose LLMEdgeRefine an iterative mechanism that forms super-points to mitigate outliers and reassign ambiguous edge points, resulting in clusters with higher coherence and robustness.
@@ 9 Discussion
We release data and code to reproduce all results in this paper at https://github.com/jairoadiazr/summaryCentroids (see the README for instructions).
Acknowledgments
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) under grant DGECR-2022-04531. The author thanks Mumin Jia, the anonymous reviewers, and the session chair for their valuable feedback and insightful suggestions, which greatly improved the quality of this work.
@@ A.1 Non-streaming datasets
We evaluate our clustering approach on four benchmark datasets. Bank77 (Casanueva et al., 2020): Consists of 3,080 customer queries related to banking services, categorized into 77 distinct intents. CLINC (Larson et al., 2019): A diverse set of 4,500 queries spanning 150 intent classes across multiple domains, designed for open-domain intent classification. GoEmo (Demszky et al., 2020): Contains 2,984 social media posts annotated with 27 fine-grained emotion categories. We removed the neutral expressions to address data imbalance and retained only entries with a single, unique emotion. MASSIVE (FitzGerald et al., 2023): Comprises 2,974 English-language virtual assistant utterances grouped into 18 domains and 59 intent categories.
@@ A.2 New compiled dataset for testing text-Streaming Clustering Algorithms
We extract and unify a challenging data stream comprising unique archive posts collected from 84 Stack Exchange sites (StackExchange, 2024). Each post is accompanied by the site label (domain) and timestamp, making this dataset well-suited for evaluating online or sequential clustering methods. Our raw dataset spans 84 domains, each containing at least 20 posts per year from 2018 to 2023 (with post lengths ranging from 20 to 1000 characters), totaling 499,359 posts. For our experiments, we focus on posts from 2020 to 2023 and further filter out labels that do not exceed 500 posts in 2023. The resulting subset comprises 35 distinct groups and 69,147 posts. Both the raw and clean data are provided with this paper. Stack Exchange content is licensed under the Creative Commons Attribution-ShareAlike 4.0 International (CC BY-SA 4.0) license.
@@ Appendix D Algorithms
    $\{\bm{\mu}_{1},\dots,\bm{\mu}_{k}\}\leftarrow\texttt{k-means++}(\{d_{1},\dots,d_{n}\},k)$;
@@
       $\bm{\mu}_{j}\leftarrow\text{Embedding}\left(\text{j}^{\text{th}}\text{ cluster summary}\right)$;
@@
       $\bm{\mu}_{j}\leftarrow\frac{1}{|C_{j}|}\sum_{i\in[C_{j}]}\mathbf{x}_{i}$;
@@
return $\{\bm{\mu}_{1},\dots,\bm{\mu}_{k}\},\{s_{1},\dots,s_{k}\}$
@@
$\{\bm{\mu}_{1},\dots,\bm{\mu}_{k}\}\leftarrow\{{\bf 0},\dots,{\bf 0}\}$;
@@
 $\{\bm{\mu}^{*}_{1},\dots,\bm{\mu}^{*}_{k}\},\{C^{*}_{1},\dots,C^{*}_{k}\},S_{b}\leftarrow\texttt{k-NLPmeans}(D_{i},k,I,m,l,T)\text{ or }\texttt{k-LLMmeans}(D_{i},k,I,m,l,T)$;
@@
return $\{\bm{\mu}_{1},\dots,\bm{\mu}_{k}\},\{S_{1},\dots,S_{b}\}$